In a clinical trial, 19 out of 800 patients taking a prescription drug complained of flulike symptoms. Suppose that it is known that 1.7% of the patients taking competing drugs complain of flulike symptoms. Is there sufficient evidence to conclude that more than 1.7% of this drug's users experience flulike symptoms as a side effect at the a=0.05 level of significance?
(a) what are the null and alternative hypothesis?
(b) Use technology to find the p-valve
Please help with this. I am struggling with this course. I need to see the work to understand what I am doing. The instructor is not much help. Thanks!

Answer by ewatrrr(24785)

Ho: p = 0.017
Ha: p > .017
-----------------------------------------
sample proportion = p-hat = 19/800 = 0.02375
---
test stat:: z(0.02375) = (.024-.017)/sqrt[0.017*0.983/800] = .007/.00457 = 1.53
-----
p-value = P(z > 1.53) = 1 - P(z<1.53) = 1 - .94 = .06
---------------------
Conclusiion:
Since the p-value is greater than 5%, accept Ho.
there is not sufficient evidence to conclude that more than 1.7% results